OpenStudy (bluebeta):

h(x)=x^2/(x^2+4) what is f and what is g?

OpenStudy (bluebeta):

@RCCB

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

are you supposed to write it as \[f(g(x))\] for some \(f\) and \(g\)?

OpenStudy (bluebeta):

@misty1212 the question says express the following function, h(x) as a composition of two functions f and g.

OpenStudy (bluebeta):

would you like me to post a picture of the problem? @misty1212

OpenStudy (misty1212):

you have a ton of choices

OpenStudy (misty1212):

perhaps the easiest one is to say \(g(x)=x^2\) and \(f(x)=\frac{x}{x+4}\)

OpenStudy (misty1212):

that way, \[f(g(x))=f(x^2)=\frac{x^2}{x^2+4}\]

OpenStudy (bluebeta):

so g(x)=x^2 and f(x)=x/x+4 ?? @misty1212

OpenStudy (bluebeta):

also i have one more question @misty1212

OpenStudy (misty1212):

ok

OpenStudy (bluebeta):

22. A ball is thrown upward with an initial velocity of 4 ft/sec from a height of 10 feet. What is the maximum height that the ball will reach?

OpenStudy (bluebeta):

and would I write both things you put for my first question?

OpenStudy (misty1212):

\[h=-16t^2+4t+10\] max is at \[-\frac{b}{2a}=-\frac{4}{-32}=\frac{1}{8}\] plug in \(f\rac{1}{8}\) to get your answer

OpenStudy (misty1212):

oops plug in \(\frac{1}{8}\) to get your answer

OpenStudy (bluebeta):

what's \rac??

OpenStudy (bluebeta):

plug in 1/8 to where? v,t, or s?

OpenStudy (misty1212):

\(t\) in \(-16t^2+4t+10\)

OpenStudy (misty1212):

looks like you get \(\frac{41}{4}\) http://www.wolframalpha.com/input/?i=-16t^2%2B4t%2B10

OpenStudy (bluebeta):

@misty1212 so do i write t=41/4?